The following refers to the hypothetical open economy where prices,wage rates and the rate of interest are constant
C=1200+0.47yd ,I=500 ,T=0.15y
Find the equilibruim level of national income ?
What is the investment multipler associated with the question ?
What is govenment multiplier associated with the question ?
By how much should government increase spending f they want the national income level to increase to 8000?
By how much should government decrease taxes if they want national income level to be 8000?
1
Expert's answer
2019-04-26T10:57:10-0400

C=1200+0.47yd, I=500, T=0.15y.

The equilibruim level of national income is:

Y = C + I = 1200 + 0.47×(Y - 0.15Y) + 500,

0.6Y = 1700,

Y = 2830.

The investment multipler associated with the question is 1.

The govenment multiplier associated with the question is 1/0.15 = 6.67.

The government should increase spending by 8000 - 2830 = 5170 if it wants the national income level to increase to 8000.

The government should decrease taxes by 5170/6.67 = 755 if they want national income level to be 8000.